OpenStudy (acespeedfighter):

5x + 1/2 y2 What is the value of the expression above when x = 3 and y = 2? You must show all work and calculations to receive full credit. (4 points)

OpenStudy (luffingsails):

Hey, I assume that y2 is "y squared"? Have you tried replacing x with 3 and y with 2 in the equation and seeing what you get?

OpenStudy (acespeedfighter):

i dont understand what your trying to say

OpenStudy (luffingsails):

Well, think about this... if you have the equation x+y and I said x =2 and y = 3. What is the value of x+y? You'd say x + y = 2 + 3 = 5. Same applies with your problem.

OpenStudy (acespeedfighter):

ohhhi see

OpenStudy (rolldextro):

He means that his problem is this \[5x+\frac{1}{2}*y^2\]

OpenStudy (luffingsails):

Yep, just plug and chug... 5(3) + 1/2 * (2)^2 = 15 + 1/2 * 4 = 15 + 2 = 17.

OpenStudy (acespeedfighter):

so the answer is 17?

OpenStudy (acespeedfighter):

oh i see

OpenStudy (luffingsails):

Right, but don't forget the work for full credit.

OpenStudy (acespeedfighter):

i know thx guys

OpenStudy (luffingsails):

You're welcome.